A baseball team has home games on Friday

and Sunday
.
The two games together earn ​$3268.50

for the team. Friday
​'s
game generates ​$603.50

less than Sunday
​'s
game. How much money was taken in at each​ game?

Found 3 solutions by mananth, TeachMath, ikleyn:
Answer by mananth(16949)

A baseball team has home games on Friday
and Sunday
.
The two games together earn ​$3268.50
for the team. Friday
​'s
game generates ​$603.50
less than Sunday
​'s
game. How much money was taken in at each​ game?
Let Sunday ​'s game generate $ x
Fridays's game generates x - 603.50
The two games generate 3268.50
x + (x-603.50) = 3268.50
2x - 603.5= 368.50
2x = 6238.50 - 603.50
2x = 5635
x= 5636/2
x = 2816
Sunday $ 2816
you may find the other





Answer by TeachMath(96)

Sunday: $1,936
Friday: $1,332.5

Answer by ikleyn(53979)

.
A baseball team has home games on Friday and Sunday
The two games together earn $3268.50 for the team.
Friday's game generates $603.50 less than Sunday's game.
How much money was taken in at each game?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth contains several arithmetic errors that lead to wrong answer.
        So, I came to bring a correct accurate solution.


Let Sunday's game generate $ x
Friday's game generates x - 603.50  dollars.

The two games generate 3268.50

    x + (x-603.50) = 3268.50

    2x - 603.5= 3268.50

    2x = 3268.50 + 603.50

    2x = 3872

     x = 3872/2

     x = 1936 

ANSWER.  Sunday's game generated  $1936.  Friday's game generated  $1936 - $603.50 = $1332.50.

Solved correctly.